# Imaging heat transport in suspended diamond nanostructures with integrated spin defect thermometers

### v2

@@ Acknowledgements.
- Dragašević and Simoncelli [2024] J. Dragašević and M. Simoncelli, Viscous heat backflow and temperature resonances in extreme thermal conductors (2024), arXiv:2303.12777 [cond-mat.mtrl-sci] .
@@
- Bayn et al. [2011] I. Bayn, A. Bolker, C. Cytermann, B. Meyler, V. Richter, J. Salzman, and R. Kalish, Diamond processing by focused ion beam—surface damage and recovery, Applied Physics Letters 99, 183109 (2011).
- Katcho et al. [2014] N. A. Katcho, J. Carrete, W. Li, and N. Mingo, Effect of nitrogen and vacancy defects on the thermal conductivity of diamond: An ab initio green’s function approach, Phys. Rev. B 90, 094117 (2014).
@@ I.2 Experimental setup
All experiments are carried out in our home-built confocal microscope. The diamond sample is mounted on a temperature-controlled stage, kept at a constant temperature of 22 ∘C for experiments in Fig. 2 and Fig. 3. The output beam of a 515 nm laser is split into two beams, one for heating the chromium patch at the apex of the cantilever and the other one for reading out the NV centers’ fluorescence all along the cantilever. The power of each beam can be controlled independently. The two beams are focused on different positions of the cantilevers with a high numerical aperture microscope objective (NA=0.9). The fluorescence is collected into a multi-mode fiber and detected with a multi-pixel photon counter (Hamamatsu C14455-1550GA).
Microwaves are delivered to the sample with a wire loop placed in close vicinity to the diamond sample. The microwave frequency is sinusoidally modulated at 5 kHz by the microwave generator (Rhodes&Schwarz SMB 100A) to be used in combination with the lock-in amplifier. The readout parameters are chosen to optimize temperature sensitivity: the readout laser power is 0.2 mW, the microwave power is 41 dBm and the microwave frequency modulation depth is 5 MHz (corresponding to approximately half of the ODMR linewidth). We have checked that the above parameters do not affect the measured value of the zero-field splitting (ZFS), confirming that the spin readout induces negligible local heating inside the cantilever. We have also verified that the heating laser does not induce any change in ZFS when focused on other portions of the cantilever, away from the metal patch.
@@ I.3 Independent study of the patch temperature
The relationship between and the lattice temperature is in general not straightforward, specifically under continuous wave illumination [77]. Hence, we perform a calibration measurement to relate the extracted to the patch temperature, using the temperature-controlled stage on which the diamond sample is mounted (see inset of Fig. S1b). We find that the relation between and the controlled patch temperature can be well described using a linear fit: . We can therefore determine the temperature of metal patches versus heating laser power, on cantilevers with different widths, as shown in Fig. S1b. Although this method does not grant high precision (note also that the extracted zero-laser-power temperature is significantly lower on the 1.18 µm cantilever), this measurement proves that the metal patch temperature increases linearly when heating laser is applied. Temperature estimates with anti-Stokes emission are also found to be of the same order as values presented in the main text on diamond cantilevers for similar heating power, meaning that the interface thermal resistance between the patch and diamond is sufficiently small. Our main conclusion is that convective and radiative thermal losses from the metal patch to the surrounding can safely be neglected.
@@ I.4 Experimental estimate of the effective thermal conductivity
The effective thermal conductivity of our cantilevers is defined as . The temperature gradient is directly extracted from the measured temperature profile. We only need to estimate the heat flux density in order to obtain . As introduced in the main text, we have . The heating laser power incident on the Cr patch is measured directly. The Cr patch absorption coefficient is estimated by measuring its reflection coefficient and assuming negligible transmission and scattering, i.e. . Finally, the cantilever cross-section is , where both the width and cross-section angle are estimated from SEM images of the different cantilevers, see for example Fig. 3a of the main text.
We assume an uncertainty of 50 nm in the extracted width from SEM images. This uncertainty needs to be properly taken into account in the uncertainty of , together with the uncertainty on the temperature gradient. We find that the width uncertainty is the dominant factor, explaining the significantly bigger error bars in , see Fig. 4, compared to the temperature gradient in Fig. 3c.
@@ I.6 Theoretical analysis
where is phonon distribution (, with being the phonon wavevector and the phonon mode index). The LBTE is a microscopic model, which describes the propagation and scattering of phonons in the presence of a temperature gradient. The LBTE can be parametrized from first principles, and its linear-response solution allows to predict the thermal conductivity. We note, in passing, that the LBTE has been recently generalized to account not only for the ability of phonons to propagate particle-like, but also for their capability to tunnel wave-like [81, 82, 83]. In diamond, propagation dominates over tunneling and the LBTE accurately describes thermal transport [14].
The first-principles solution of the LBTE has the drawback of being computationally expensive, making it unpractical to explore how nanostructure shape affects heat transport. Therefore, to investigate thermal transport in cantilevers having different cross-section shapes and sizes, we rely on the mesoscopic viscous heat equations (VHE) [14]. The VHE are two coupled partial differential equations for the mesoscopic temperature () and phonon drift-velocity () fields. They are obtained from coarse-graining the LBTE[84, 14]. The VTE encompass both the diffusive (Fourier) and hydrodynamic regimes [7, 84] of thermal transport, as well as the crossover between them. It has been shown that the VHE rationalizes the experimental demonstration of hydrodynamic transport in graphite [14], and also yields predictions compatible with the space-dependent solution of the LBTE[57, 58] in complex-shaped devices [59]. In the steady-state regime, for a material with isotropic transport properties such as diamond, the VHE read:
@@
In the limit of negligible hydrodynamic effects, Eqs. (S3) reduce to the Fourier’s law of diffusive heat conduction:
In this regime the heat flux simply becomes
The importance of viscous effects can thus be estimated by comparing the results of computing the heat flux with Eq. (S4) (full solution) versus Eq. (S7) (purely diffusive case).
@@
All the parameters entering the VHE can be determined from the first-principles solution of the full LBTE (i.e. accounting for the actual phonon band structure and full collision matrix). An exact solution of the LBTE has been discussed for the case of nanowires [49, 60] under the assumption of diffusive boundaries. This approach allows to account for the influence of boundary scattering on the conductivity of wires, providing an expression which depends only on the cross-section of the wire 111In particular, Eq. 6 of Ref. [49] encompasses as special cases both thin films and thin wires, as discussed after Eq. 5 of Ref. [67] and Eq. 11.2.9 of Ref. [68] and can be applied to the triangular cross-section nanowire studied in this work. After verifying with the established Casimir model [50] that boundary scattering has strong effects on the conductivity but has negligible effect on the other parameters entering the VHE (including the thermal viscosity ), we accounted for nanowire boundary scattering on the conductivity only and used bulk values (calculated in Ref. [14]) for the other parameters entering in the VHE. This approximation is further justified a posteriori, since Fig. 4 shows that viscous effects are unimportant in the cases considered here.
Finally, we discuss how we considered the Ga contamination in the vicinity of milled faces of the cantilevers (i.e. bottom faces, see inset of Fig. 4). Since with the methods currently available it is not possible to solve the LBTE with position-dependent Ga content, we consider uniform Ga impurities over the whole cantilever cross-section. This approximation allows to capture the increased relative influence of the Ga-contaminated layer (of fixed depth , see main text) when reducing the cantilever width . In particular, we consider an impurity concentration proportional to the ratio of contaminated area to total cross-section area, referred to as penetration ratio
where the approximated equality holds at linear order in ; details on the geometrical analysis performed to compute PR are reported in the next section. Eq. (S8) allows us to describe the width-dependent average Ga concentration as:
where is the concentration of Ga impurities in the contaminated layer. We thus determine the conductivity from the linear-response solution of the LBTE, accounting for phonon-impurity scattering with impurity strength parameter corresponding to Ga impurities with concentration [61].
@@
In summary, we determine the linear-response solution of the LBTE (Eq. (S2)) for diamond nanowires with width-dependent Ga impurity scattering using the ShengBTE package [60] and use the bulk thermal viscosity computed following the theoretical prescriptions of Ref. [14]. Then, we solve the VHE numerically (Eq. (S3)) using a finite-element solver implemented in Mathematica [86], imposing a temperature gradient K around room temperature on µm long cantilevers. Finally, we obtain the effective thermal conductivity from Eqs (S4) and (S5). The Ga concentration in the contaminated areas is set in the ShengBTE simulation so as to fit the experimental trend of vs width (black squares in Fig. 4 of the main text). This is the sole fitting parameter in our model.
I.7 Evaluation of the Ga participation ratio
Given that Ga impurities are mostly localized on the two slanted sides of Diamond cantilevers (see inset of Fig. 4), we can compute the participation ratio, PR, as the ratio between the azure area in the triangle on the right of Fig. 4, , to the total triangular section, . Considering an angle at the surface equal to , with some algebra one finds
where the fraction multiplied by a factor of 2 represents each of the -deep areas related to the two slanted sides (see the inset of Fig. 4 of the main text). On the other hand, the total triangular section is simply
These considerations allow us to approximate to linear order in and justify Eq. (S8) as shown below:
The solution of the LBTE with a space-dependent scattering term for nanowires is obtained using the ShengBTE package. Even if the solution is valid for any cross-section shape (it depends only on the cross-section value), the ShengBTE implementation is coded assuming a circular cross-section. So, the widths of the prismatic cantilevers are obtained from the ShengBTE simulations by matching circular cross-section used in the simulation, ( being the radius of the cylindrical nanowire simulated), with the triangular cross-section, . This simply translates into getting the width values of triangular nanowires (used on the x axis of Fig. 4 of the main text) as .
Finally, in Fig. S7, we show further effective thermal conductivity calculations obtained using a fixed (non width-resolved) Ga-impurity concentration. We clearly observe that the slope of such curves does not capture the experimental trend, highlighting the necessity to consider a Ga-impurity model.

### v3

@@ Acknowledgements.
- Aharon-Steinberg et al. [2022] A. Aharon-Steinberg, T. Völkl, A. Kaplan, A. K. Pariari, I. Roy, T. Holder, Y. Wolf, A. Y. Meltzer, Y. Myasoedov, M. E. Huber, et al., Direct observation of vortices in an electron fluid, Nature 607, 74 (2022).
- Palm et al. [2024] M. L. Palm, C. Ding, W. S. Huxter, T. Taniguchi, K. Watanabe, and C. L. Degen, Observation of current whirlpools in graphene at room temperature, Science 384, 465 (2024).
@@
- Gomès et al. [2015] S. Gomès, A. Assy, and P.-O. Chapuis, Scanning thermal microscopy: A review, physica status solidi (a) 212, 477 (2015).
- Zhang et al. [2020] Y. Zhang, W. Zhu, F. Hui, M. Lanza, T. Borca-Tasciuc, and M. Muñoz Rojo, A review on principles and applications of scanning thermal microscopy (sthm), Advanced Functional Materials 30, 1900892 (2020).
@@
- Laraoui et al. [2015] A. Laraoui, H. Aycock-Rizzo, Y. Gao, X. Lu, E. Riedo, and C. A. Meriles, Imaging thermal conductivity with nanoscale resolution using a scanning spin probe, Nature Communications 6, 8954 (2015).
@@
- Dragašević and Simoncelli [2024a] J. Dragašević and M. Simoncelli, Viscous heat backflow and temperature resonances in extreme thermal conductors (2024a), arXiv:2303.12777 [cond-mat].
@@
- Carruthers [1961] P. Carruthers, Theory of thermal conductivity of solids at low temperatures, Reviews of Modern Physics 33, 92 (1961).
- Note [1] We note that this is formally analogous to Bosanquet-type regularizations for the transport coefficients of rarefied fluids [75, 62].
- Cepellotti and Marzari [2017] A. Cepellotti and N. Marzari, Boltzmann transport in nanostructures as a friction effect, Nano Letters 17, 4675 (2017).
- Maassen and Lundstrom [2015] J. Maassen and M. Lundstrom, Steady-state heat transport: Ballistic-to-diffusive with fourier’s law, Journal of Applied Physics 117 (2015).
@@
- Michalis et al. [2010] V. K. Michalis, A. N. Kalarakis, E. D. Skouras, and V. N. Burganos, Rarefaction effects on gas viscosity in the Knudsen transition regime, Microfluidics and Nanofluidics 9, 847 (2010).
@@
- Dragašević and Simoncelli [2024b] J. Dragašević and M. Simoncelli, Viscous heat backflow and temperature resonances in extreme thermal conductors (2024b), arXiv:2303.12777 [cond-mat.mtrl-sci] .
@@
- Dragašević and Simoncelli [2023] J. Dragašević and M. Simoncelli, Viscous heat backflow and temperature resonances in extreme thermal conductors, arXiv preprint arXiv:2303.12777 (2023).
@@
- Allen and Perebeinos [2018] P. B. Allen and V. Perebeinos, Temperature in a peierls-boltzmann treatment of nonlocal phonon heat transport, Physical Review B 98, 085427 (2018).
- Sendra et al. [2022] L. Sendra, A. Beardo, J. Bafaluy, P. Torres, F. X. Alvarez, and J. Camacho, Hydrodynamic heat transport in dielectric crystals in the collective limit and the drifting/driftless velocity conundrum, Physical Review B 106, 155301 (2022).
- Gürcan et al. [2013] Ö. D. Gürcan, P. Diamond, X. Garbet, V. Berionni, G. Dif-Pradalier, P. Hennequin, P. Morel, Y. Kosuga, and L. Vermare, Transport of radial heat flux and second sound in fusion plasmas, Physics of Plasmas 20 (2013).
@@ I.2 Experimental setup
All experiments are carried out in our home-built confocal microscope. The diamond sample is mounted on a temperature-controlled stage, kept at a constant temperature of 22 ∘C (except for calibration measurements in Fig. 1). The output beam of a 515 nm laser is split into two beams, one for heating the chromium patch at the apex of the cantilever and the other one for reading out the NV centers’ fluorescence all along the cantilever. The power of each beam can be controlled independently. The two beams are focused on different positions of the cantilevers with a high numerical aperture microscope objective (NA=0.9). The fluorescence is collected into a multi-mode fiber and detected with a multi-pixel photon counter (Hamamatsu C14455-1550GA).
Microwaves are delivered to the sample with a wire loop placed in close vicinity to the diamond sample. The microwave frequency is sinusoidally modulated at 5 kHz by the microwave generator (Rohde&Schwarz SMB 100A) to be used in combination with the lock-in amplifier. The readout parameters are chosen to optimize temperature sensitivity: the readout laser power is 0.15 mW, the microwave power is 41 dBm and the microwave frequency modulation depth is 5 MHz (corresponding to approximately half of the ODMR linewidth).
Fig. S2 shows how the above parameters affect the measured value of the zero-field splitting (ZFS). We find that, on cantilevers fabricated with FIB milling, the readout laser induces a small heating (blue squares). On the other hand, no laser-induced heating is measured on cantilevers fabricated only with Faraday cage angle etch, without the final FIB milling step (red triangles). We conclude that the layer contaminated with Ga ions during milling causes absorption of the readout laser light, which in turn generates local heating. Nevertheless, the small, K, heating caused by the readout laser with 0.15 mW power is not expected to affect the temperature profile measured in the presence of the heating laser on the metal patch: this contribution from the readout laser is removed when subtracting the ZFS values without the heating laser (see Fig. S8). The microwave power is found to have a negligible effect on the temperature readout. Note that no static magnetic field is applied to the sample. This magnetic field-free scheme to measure the ZFS for thermometry is known to be insensitive to small magnetic field fluctuations [51].
@@ I.3 Independent study of the patch temperature
The relationship between and the lattice temperature is in general not straightforward, specifically under continuous wave illumination [85]. Hence, we perform a calibration measurement to relate the extracted to the patch temperature, using the temperature-controlled stage on which the diamond sample is mounted (see inset of Fig. S3b). We find that the relation between and the controlled patch temperature can be well described using a linear fit: . We can therefore determine the temperature of metal patches versus heating laser power, on cantilevers with different widths, as shown in Fig. S3b. Although this method does not grant high precision (note also that the extracted zero-laser-power temperature is significantly lower on the 1.18 µm cantilever), this measurement proves that the metal patch temperature increases linearly when heating laser is applied. Temperature estimates with anti-Stokes emission are also found to be of the same order as values presented in the main text on diamond cantilevers for similar heating power, meaning that the interface thermal resistance between the patch and diamond is sufficiently small.
We can estimate convective and radiative thermal losses from the metal patch using the above temperature measurements. For the range of applied laser power, we get patch temperature . We use this value to get an upper-bound estimate of radiative and convective losses. For radiative heat transfer, we use the Stefan-Boltzmann law for black-body radiation:
where is the Stefan-Boltzmann constant and is the emissivity of the body. Even for a perfectly black body, i.e., , we get an upper bound W. Convective heat transfer can be estimated as
with the patch area and the ambient air temperature. Considering a value of the heat transfer coefficient for free convection with air (which is on the higher end of values reported in the literature), µm2 and , we get an estimated W. Both estimates for and are 5 to 6 orders of magnitude lower than the applied heating laser power. Our main conclusion is thus that convective and radiative thermal losses from the metal patch to the surrounding can safely be neglected. The same is also true for radiative and convective losses in the diamond cantilever itself (it has even lower temperature and its outer surface is only 10 to 100x bigger than the patch).
@@ I.4 Experimental estimate of the effective thermal conductivity
The effective thermal conductivity of our cantilevers is defined as . The temperature gradient is directly extracted from the measured temperature profile. We only need to estimate the heat flux density in order to obtain . As introduced in the main text, we have . The heating laser power incident on the Cr patch is measured directly. The Cr patch absorption coefficient is estimated by measuring its reflection coefficient and assuming negligible transmission and scattering, i.e. . Finally, the cantilever cross-section is , where both the width and cross-section angle are estimated from SEM images of the different cantilevers, see for example Fig. 1b and Fig. 3a of the main text. Note that the explicit expression for as a function of the above experimental parameters, specifically and , is:
We conservatively assume an uncertainty of 50 nm in the extracted width from SEM images. This uncertainty needs to be properly taken into account in the uncertainty of , together with the uncertainty on the temperature gradient. We find that the width uncertainty is the dominant factor, explaining the significantly bigger error bars in , see Fig. 4, compared to the temperature gradient in Fig. 3c.
@@ I.6 Theoretical analysis
where is phonon distribution (, with being the phonon wavevector and the phonon mode index). The LBTE is a microscopic model, which describes the propagation and scattering of phonons in the presence of a temperature gradient. The LBTE can be parametrized from first principles, and its linear-response solution allows to predict the thermal conductivity. We note, in passing, that the LBTE has been recently generalized to account not only for the ability of phonons to propagate particle-like, but also for their capability to tunnel wave-like [89, 90, 91]. In diamond, propagation dominates over tunneling and the LBTE accurately describes thermal transport [61]. Moreover, since in diamond momentum-conserving “normal” phonon scattering processes have higher (or comparable) strength than momentum-non-conserving ”Umklapp” processes, a solution of the LBTE accounting for the full scattering matrix is necessary to properly capture heat-transport properties [53, 54].
The first-principles solution of the LBTE with full scattering matrix has the drawback of being computationally expensive, making it impractical to explore how nanostructure shape affects heat transport. Therefore, to investigate thermal transport in cantilevers having different cross-section shapes and sizes, we rely on the mesoscopic viscous heat equations (VHE) [61]. The VHE are two coupled partial differential equations for the mesoscopic temperature () and phonon drift-velocity () fields. They are obtained from coarse-graining the LBTE[92, 61]. The VTE encompasses both the diffusive (Fourier) and hydrodynamic regimes [7, 92] of thermal transport, as well as the crossover between them. It has been shown that the VHE rationalizes the experimental demonstration of hydrodynamic transport in graphite [61], and also yields predictions compatible with the space-dependent solution of the LBTE[64, 65] in complex-shaped devices [93]. In the steady-state regime, for a material with isotropic transport properties such as diamond, the VHE read:
@@
All the parameters entering the VHE can be determined from the first-principles solution of the full LBTE (i.e. accounting for the actual phonon band structure and full collision matrix). An exact solution of the LBTE has been discussed for the case of nanowires [53, 66] under the assumption of diffusive boundaries. This approach allows to account for the influence of boundary scattering on the conductivity of wires, providing an expression which depends only on the cross-section of the wire 222In particular, Eq. 6 of Ref. [53] encompasses as special cases both thin films and thin wires, as discussed after Eq. 5 of Ref. [76] and Eq. 11.2.9 of Ref. [60] and can be applied to the triangular cross-section nanowire studied in this work. Importantly, the description of finite-size effects is based on the characteristic length scale of the problem, following Casimir’s treatment of boundary scattering in small devices at low temperatures [68]. In this regime, where intrinsic anharmonic phonon scattering occurs at a rate comparable to or higher than boundary scattering, the appropriate length scale to use is one-half of the smallest system dimension [70]. After verifying with the established Casimir model [54] that boundary scattering has strong effects on the conductivity but has negligible effect on the other parameters entering the VHE (including the thermal viscosity ), we accounted for nanowire boundary scattering on the conductivity only and used bulk values (calculated in Ref. [61]) for the other parameters entering in the VHE. This approximation is further justified a posteriori, since Fig. S10 shows that viscous effects are unimportant in the cases considered here.
@@
In summary, we determine the linear-response solution of the LBTE (Eq. (S5)) for diamond nanowires using the ShengBTE package [66] and use the bulk thermal viscosity computed following the theoretical prescriptions of Ref. [95]. Then, we solve the VHE numerically (Eq. (S6)) using a finite-element solver implemented in Mathematica [96], imposing a temperature gradient K around room temperature on µm long cantilevers. Finally, we obtain the effective thermal conductivity from Eqs (S7) and (S8). There is no fitting parameter in our model.
I.7 Estimation of the magnitude of viscous hydrodynamic effects
Violations from standard diffusive thermal transport can emerge in multiple ways: when microscopic carriers achieve a mean free path comparable to the device size [97], or when momentum-conserving phonon-phonon collisions dominate over momentum-relaxing Umklapp collisions [61]. Both ballistic and hydrodynamic effects share the common underpinnings of being described by non-local relations between temperature gradient and heat flux [98, 99, 95]. The problem of distinguishing these effects is challenging and not addressed even in pioneering works, which employ phenomenologically parametrized mesoscopic models [14, 15] or simplified microscopic models that do not capture hydrodynamic physics [30]. Depending on the relative magnitude of hydrodynamic and size effects, we can distinguish different regimes [71]: (i) “ideal diffusion”, where the thermal conductivity is an intrinsic constant; (ii) “apparent diffusion”, the regime in which anomalous, non-Fourier size-heat flux relations are driven by a conductivity that depends on size due to ballistic effects; (iii) “hydrodynamics”, which originates from the presence of viscous effects dominating over momentum-dissipation effects.
To distinguish between these different regimes, we estimate the magnitude of hydrodynamic effects in the cantilevers. In the limit of negligible hydrodynamic effects, Eqs. (S6) reduce to :
In the absence of ballistic effects, Eq. S9 is exactly Fourier’s law of diffusive heat conduction. In the presence of ballistic effects, the thermal conductivity can no longer be considered an intrinsic material property and depends on size, corresponding to the apparent diffusive regime. In this regime, the heat flux becomes
The importance of viscous effects can thus be estimated by comparing the results of computing the heat flux with Eq. (S7) (full solution) versus Eq. (S10) (apparent diffusive case). Fig. S10 shows the effective thermal conductivity computed in both cases, together with experimental data presented in the main text. The two curves show a minor difference over the whole range of width. This means that viscous effects only play a minor role in the observed deviations from pure diffusion. Ballistic effects are thus the main factor in the reduction of for reducing in 1-dimensional diamond cantilevers.
I.8 Influence of vacancy impurities